40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog  Not logged in ELOG logo
Message ID: 975     Entry time: Wed May 13 17:10:16 2015
Author: Gabriele 
Type: DailyProgress 
Category:  
Subject: Some retuning 
  • The Michelson fringes looked too fast to me, so I started investigating the reason. The z damping od both test blades was not very efficient: it reduced the motions but noth really overdamping the blades. So I retuned the loop shape, moving the elliptic low pass to a slightly higher frequency and increasing the gain by about a factor 5. The 'notch' filter bank is no more necessary, since it was used to reduce the correction by killing some lines in the optical table motion, which are no more there. 
    With this new configuration of the damping, the 2.2 Hz resonances are overdamped, and fringes are significantly slower
  • The locking filter we designed for the lock on tha table was working, even though with not good performances. I modified it, moving down the roll-off poles to about 10 Hz. I also modified the elliptic low pass, moving it to 20 Hz and adding an additional low pass at 100 Hz. In this way I reduced a lot the high frequency control signal. Still, I'm not able to lock with very good performance: the error signal is sitting around 0, but with a peak to peak variation of the order of half or full fringe. My impression is that we have a lot of motion due to air currents, but this has to be verified.
ELOG V3.1.3-